12/18/06 – Final payment and pay application has been received from Stapleton. TEI is forwarding to Kathy Buschbacher for closeout. This packet does include closeout documents.

11/30/06 – TTC has been working with Stapleton to get closeout complete and the final security key switch complete. It appears the job is ready for closeout. TTC contacted Langdon about closeout documents and payroll today as well. Langdon reported that the final balance report just arrive last week and that O&M’s are probably being completed right now. Kathy has asked for closeout on this project and TEI is trying to get that done in the next 2 weeks. TTC emailed CRA about status of Langdon closeout as well along with allowance usage plan for Stapleton. Allowance Usages for Stapleton: (1) $677.05 – Change Order # 01 – worked on wireless pushbuttons (stapleton bills directly to zeiser for payment from the GC allowance) (2) $410.05 – Change Order # 02 – Install Transmitter, Receiver, Transformer (from stapleton allowance) (3) $567.06 – Change Order # 03 – Access Panel in Ceiling, Fish Door Hardware, Cables to Controllers (from stapleton allowance) (4) $1437.00 – Change Order # 04 – Wireless RF Transmitters with relay and box, Wire door opener, temporary switches for control (stapleton bills directly to zeiser for payment from the GC allowance) (5) $409.85 – Change Order # 05 – Remove Controller, Wire door controls, add relay & transmitter (stapleton bills directly to zeiser for payment from the GC allowance) (6) $598 – Change Order # 06 – Key switch for Security Override (stapleton bills directly to zeiser for payment from the GC allowance)

11/17/06 – TE has performed the punchlist for this project. There are items left on the control of the door hardware for Stapleton to take care of but that seems to be a problem to get finished at this point due to the precarious nature of the control wiring to have the doors work from the standpost and by wireless and then to be able to disable them at night when the building needs to be secured. CRA and TE have put Stapleton in touch with Ingersoll Rand to facilitate the final connections. 11/17/06 was to be the final date this work was to get done. TE approved Stapleton’s #1 and #2 pay applications today and had them City Dashed down to the A&D building for payment.

07/25/06 – Const Mtg today revealed that we have redesigned the door hardware issues that are associated with project. Very little progress has been made except that sprinklers are in, the diffusers are being installed and six lights have been installed. The door hardware issues need to be cleared up so the rest of the job can move forward.

07/14/06 – The coordination and work over the last two weeks includes modifying the plenum such that a continuous sheet metal wall was installed for the plenum air for the boiler room and elbows were added from the existing louvers to the face of the this wall for proper airflow. The ceiling was raised in the space per Dave Woodside desires and this created a clearance problem for the new can lights being installed because our ceiling was too close to the our new ductwork. Mike Stiles resolved these issues with the electrician by relocating the lights into a horseshoe pattern but this also means that the ceiling contractor has to move some of the new studs to make this work. TTC approved Langdon’s #2 invoice on 7-14-06 for payment. Mike and Spencer found a method to resolved the handicap access and power to the new door actuator’s.

05/24/06 – TTC talked to Bill Bauer of Langdon today. They believe that TEI’s new layout is acceptable and will work well in this installation. The new layout reuses the glycol piping for the fan powered box, keeps the existing louvers in place in the lobby and adds a couple of 24×24 lay-ins will provide access into the new ceiling space. All the new parts are ordered and should be here early next week. The HVAC and piping contractors should be down next Thursday to begin this work for the vestibule. I believe that would be June 1. TTC informed all partners.

05/23/06 – TTC provided a PDF file to everyone for review that makes the HVAC work much easier, re-using the glycol heating piping for the VAV box that is subject to cooler air from the return plenum. TTC contacted Langdon again today to see if they are ready to make the changes.

05/16/06 – Store front glass is partially installed. TTC is meeting Langdon on site to coordinate some ductwork above the ceiling tonight at 4:30 p.m.

04/17/06 – TEI received fire protection submittals today for review. TEI reviewed project scope with Bill Bauer at last pre const mtg. Door hardware is pushing back the project 4 weeks.

04/03/06 – TEI sent back the UH submittals today. No Exceptions Noted.

03/07/06 – TEI received submittals today on control valves for this job. TTC gave to Sam to review.

02/24/06 – Jim Lape called to inquire about mechanical submittals. TTC & SRB finished them today and city dashed them to Jim for use on this project.

01/31/06 – We had the pre construction meeting today for this job. Langdon is the MC. Stapleton is the EC for this project. Meeting Notes PDF

01/01/06 – This is a C&R lead project. TEI bid this out with C&R and then had to rebid the electric portion under a separate contract. Now that the project meetings have started we are treating the whole project as one big project with C&R as the lead and Jim Lape at Cole & Russell running the show. TTC is asking that the web journal be kept up to date on this project for record keeping purposes.